1.The core diameter of multimode step index fibre is 60 micrometer. The difference in refractive index is 0.013. The core refractive index is 1.46. Determine the number of guided modes when the operating wavelength is 0.75 micrometer.

Grade:10

1 Answers

KIPTANUI DOMINIC
2 years ago
n1 = 1.473
n2 = 1.46.
Core Diameter = 60 μm.
Wavelength = 0·75 μm
 
NA=  \sqrt{1.473^{2}}-{1.46^{2}}

N·A= 0.195
 
No. of modes =      \frac{1}{2}( \frac{pie * D}{Wavelenght} *N.A)^{2}
 
No. of modes =         \frac{1}{2}( \frac{\frac{22}{7}*D((0.75X10^{-6} ))}{(0.75X10^{-6})) } *0.195)^{2}

No. of Modes = 1200